OpenStudy (anonymous):

x = sqrt(1+i) -sqrt(1-i) for x.

OpenStudy (anonymous):

Wolfram say s the square root of i-1 and i+1 are the same...

OpenStudy (anonymous):

so zero !!! O_o I'm confused

OpenStudy (anonymous):

Can't get the link to post correctly....

OpenStudy (anonymous):

How about Polar Form ?

OpenStudy (anonymous):

Lets say z = 1 -i

OpenStudy (anonymous):

Put sqrt(1+i) - sqrt(1-i) in Wolfram....

OpenStudy (anonymous):

i sqrt (2(sqrt2-1))....

OpenStudy (anonymous):

if i do that in polar i get just 1 -i \[\sqrt{2}e^{i \frac{3\pi}{4}}\]

OpenStudy (anonymous):

then for 1+i it becomes same but with pi/4

OpenStudy (anonymous):

ANd why hasn't sqrt(1-i) or (1+i) actually got 2 roots not 1?

OpenStudy (anonymous):

I'm confused::¿

OpenStudy (anonymous):

Me too !

OpenStudy (anonymous):

Ur having same trouble with link as me, it's not right...

OpenStudy (anonymous):

Decimal Approximation .90.. something

OpenStudy (anonymous):

That what Wolfram is showing

OpenStudy (anonymous):

They must be "choosing" a square root but what basis, u can't say the "positive" one..

OpenStudy (anonymous):

Grr.. can't get the links to come up right....

OpenStudy (anonymous):

Put simplify sqrt etc in Wolfram

OpenStudy (anonymous):

Ok For Polar it should look like this \[x =\sqrt{\sqrt{2}e^{i \frac{\pi }{4}}} - \sqrt{\sqrt{2}e^{i \frac{3 \pi}{4}}} \]

OpenStudy (anonymous):

Now u have the eîpi = 1 thingy....

OpenStudy (anonymous):

So you split the pi/4 into 1 and eî whatever.

OpenStudy (anonymous):

e^(1/4)

OpenStudy (anonymous):

Yeah then you take sqrt{sqrt2 }common and it should become \[2^{1/4} \times ( e^{i \frac{\pi}{8}} - e^{i \frac{3 \pi }{8}})\]

OpenStudy (anonymous):

sqrt(sqrt2 e^(1/4))....lol

OpenStudy (anonymous):

then when we open it into e^i(theta ) into cos theta + i sin theta

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

I gave u medal for trying to help....:-)

OpenStudy (anonymous):

>What a mess!

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

with it should become \[2 \times i \times 2^{1/4} sin\frac{\pi}{8}\]

OpenStudy (anonymous):

Now we have to solve this

OpenStudy (anonymous):

Yes they have like that as alternate form in Wolfram....

OpenStudy (anonymous):

But still, they are choosing a root...

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

And they are choosing the same root for both terms!!!!

OpenStudy (anonymous):

Maybe this has something to do with it..? http://en.wikipedia.org/wiki/Branch_cut#Branch_cuts

OpenStudy (anonymous):

I am so confused now ..

OpenStudy (anonymous):

Let's leave it and come back to it later....?

OpenStudy (anonymous):

Yeah maybe others Like joemath and Zarkon can clear this ..they are really good ..I would have to read about it I'l come back later after reading enough then try this ..

OpenStudy (anonymous):

Latest question...what is the last number of the pi? :-) Let's go solve that one...

OpenStudy (phi):

\[x ^{2}=1+i+1−i+2(1−i ^{2})\] should be \[x ^{2}=1+i+1−i+2\sqrt{(1−i ^{2})}\] You will eventually get x= sqrt(stuff), so x will be plus or minus this expression. As for the \[\sqrt{i+1}\] It's most easily done in polar coordinates. e.g. \[1+i = \sqrt{2}e ^{\frac{i \pi}{4}}\] and its square root is \[(1+i)^{\frac{1}{2}}= \pm2^{\frac{1}{4}} e ^{\frac{i \pi}{8}}\] Of course you can re-write the negative root as\[2^{\frac{1}{4}}e ^{\frac{i \pi}{8}} e ^{i\pi} = 2^{\frac{1}{4}}e ^{\frac{i 9 \pi}{8}}\]

OpenStudy (anonymous):

Satellite, u got a take on this one?

OpenStudy (phi):

oops, that should be a - 2 sqrt(1- i*i) in the starting equation. i.e. \[x^{2}=1+i+1−i−2\sqrt{(1−i^{2})}\] which simplifies to \[x^{2}=2−2\sqrt{2}\] and \[x= \pm\sqrt{2(1−\sqrt{2})}\] Now, because the 1-sqrt(2) is negative, we can re-write this as \[x= \pm i\sqrt{2(\sqrt{2}-1)}\] Tough typing this in without the random mistake!

OpenStudy (anonymous):

I think I have id'd the problem with this now. It depends on whether u allow x to take complex values. The root 2 thingy is if x can take complex and the other is if x is real.

OpenStudy (anonymous):

And Wolfram seem to arbitrarily choose the expression with positive i as a kind of "positive" root and then take the other for the conjugate.

OpenStudy (phi):

Here's what I think is going on...but I haven't really checked... First, we are working on the complex plane. Now let's take the nth root of a number. The number in polar form is magnitude * exp(i theta) its first root is magnitude^(1/n) * exp(i theta/n) Of course there are n roots... we can always add 2pi to the argument (theta), so the next root is at magnitude^(1/n) * exp( i theta/n + 2*pi/n) ok let me ignore the magnitude..it's always the same... the next root is at exp(i theta/n + 4*pi/n) and so on until we have gone around the circle notice that if the number is positive real we end up at exp(i*0) and exp(ipi) (i'm ignoring magnitude) To summarize wolfram is just picking the 1st root at exp(i theta/n)

OpenStudy (phi):

*notice that if the number is positive real we end up at exp(i*0) and exp(i pi) I mean for the square root (n=2)

OpenStudy (anonymous):

I am going to go back to my geometrical interpretation of i and forget all about complex analysis...:-)

OpenStudy (phi):

I take it my explanation was not clear. Here's another try, with pictures!

hero (hero):

Oh wow....interesting